A ball is kicked upward with an initial velocity of 52 feet per second. The ball's height from the ground is modeled by h=-16t squared +52t where t is measured in seconds. How much time does the ball reach its highest point? what is the height at this point?
Answer by Alan3354(69443)

h(t) = -16t^2 +52t
---------
It's a parabola.
Find the vertex
t at the vertex = -b/2a = -52/-32
t = 1.625 seconds
--------
Max height = h(1.625)
= -16*1.625^2 + 52*1.625
= 42.25 feet